# Position: Beyond Euclidean - Foundation Models Should Embrace Non-Euclidean Geometries

### v1

@@ 3.1 Limitations of Euclidean Geometry for Foundation Models
(Robinson, 2006) Let $(X,d_{X})$ be a metric space with $X=\{a,b,c,d\}$ and $d_{X}(a,b)=d_{X}(a,c)=d_{X}(a,d)=2L$ and $d_{X}(b,d)=d_{X}(c,d)=L$ for $L\in\mathbb{R}^{+}$. Then $X$ admits no isometric embedding into $\mathbb{R}^{n}$ for any $n$.
As these points can be isometrically embedded into $S^{k}$, Theorem 3.3 shows that $S^{k}$ cannot be isometrically embedded into $\mathbb{R}^{n}$ for any $n\in\mathbb{N}$, resulting in distortion when encoding properties such as rotational equivariance. In contrast, non-Euclidean geometry (e.g., hyperbolic or spherical spaces) can provide a more natural representation of complex topological structures, reducing distortion and dimensionality of the embedding space. For instance, (Sarkar, 2011) showed that every finite tree admits an embedding into the hyperbolic plane $\mathbb{H}^{2}$ with only $1+\epsilon$ multiplicative distortion for any $\epsilon>0$. This leads to a distortion bound of $O(1)$ while requiring low dimensionality.
Table 1: Average (point-wise) distortion on canonical graphs with $96$ nodes, comparing four spaces with total dimension $6$. The least distortion is achieved by the space with the most suitable geometry.
@@
Take-away. The implications of the previous theoretical discussion are numerous: (1) Limited scalability. Theorem 3.1 highlights the distortion-dimension trade-off for Euclidean foundation models when embedding complex structures, which is reflected in the computational resources required in these models. Non-Euclidean geometry produces higher quality embeddings in significantly lower dimensions, offering enhanced model scalability; (2) Performance bottleneck. Theorem 3.2 and 3.3 demonstrate that even in the case of an abundance of compute resources, the linear assumption in Euclidean foundation models could still incur significant distortion regardless of the embedding dimension for a wide range of topological structures, resulting in a performance upper bound.
Model arXiv C4 Common Crawl GitHub StackExchange Wikipedia
RoBERTa-Base (Liu et al., 2019b) $0.15\pm 0.06$ $0.18\pm 0.04$ $0.17\pm 0.04$ $0.12\pm 0.04$ $0.17\pm 0.07$ $0.07\pm 0.05$
LLaMA3.1-8B (Grattafiori et al., 2024) $0.15\pm 0.05$ $0.16\pm 0.07$ $0.15\pm 0.06$ $0.12\pm 0.05$ $0.18\pm 0.06$ $0.10\pm 0.04$
GPT-NeoX-20B (Black et al., 2022) $0.14\pm 0.03$ $0.17\pm 0.06$ $0.15\pm 0.05$ $0.11\pm 0.04$ $0.14\pm 0.04$ $0.09\pm 0.03$
Gemma2-9B (Team et al., 2024) $0.17\pm 0.06$ $0.19\pm 0.04$ $0.20\pm 0.05$ $0.15\pm 0.05$ $0.18\pm 0.04$ $0.15\pm 0.03$
Reference $\delta$ values $0.99\pm 0.01$ $0.63\pm 0.01$ $0.40\pm 0.04$ $0.14\pm 0.01$ $0.0$
Figure 2: Token frequency v.s. token count (left 2) and token norm vs token count (right 2) for LLaMa3.1-8B and LLaMaGen. The datasets are chosen to be within the training corpus. The token-frequency figures show the scale-free properties of the token inputs. The token norms figures reflect this property for learned token embeddings to some extent, with token count decreasing exponentially for high normed tokens at the right tail. However, the Euclidean embeddings stil do not fully capture this property and deviate from it at the left tail. Statistics for additional models and dataset details are shown in Appendix B.
3.2 Non-Euclidean Geometry in Foundation Models
In this section, we empirically assess embedding distortions for different geometries to validate our claims in Section 3.1 and demonstrate that non-Euclidean geometry is more suitable. We then analyze token embeddings in foundation models, showing that structures that align with non-Euclidean geometry (Section 3.1) are prevalent, highlighting the need for alternative geometric frameworks.
@@ 3.2 Non-Euclidean Geometry in Foundation Models
Empirical Validation. We empirically validate our claim that Euclidean space fails to capture complex structures faithfully and that non-Euclidean spaces are better suited for producing high-quality embeddings. Table 1 compares the average (point-wise) distortion of four geometric spaces ($\mathbb{R}^{6}$, $\mathbb{H}^{-1,6}$, $\mathbb{S}^{1,6}$, and $\mathbb{H}^{-1,3}\times\mathbb{S}^{1,3}$) in representing three canonical graphs (Tree, Cycle, and Ring of Trees) with 96 nodes, each corresponding to a different type of intrinsic token relationships (hierarchical, cyclical, and both). The most suitable geometry varies by graph type—Lorentzian space ($\mathbb{H}^{-1,6}$) for trees, spherical space ($\mathbb{S}^{1,6}$) for cycles, and mixed geometry ($\mathbb{H}^{-1,3}\times\mathbb{S}^{1,3}$) for rings of trees—emphasizing the importance of selecting an appropriate geometry to minimize distortion.
@@
Hierarchies in Token Embeddings. To explore the intrinsic structure within the representations of foundation models, we first utilize $\delta$-hyperbolicity (Gromov, 1987)—a curvature notion that quantifies the extent to which a metric space $(X,d)$ deviates globally from a perfect tree metric. Formal details are provided in Appendix C. Each token within a prompt is treated as a point in a discrete metric space $X$, and a graph is constructed based on similarity scores between each pair. We evaluate the hyperbolicity of token embeddings in LLMs, where lower values suggest a tree-like structure (Table 3). As shown in Table 2, the consistently low $\delta$-hyperbolicity values suggest hierarchical structures within each prompt across diverse datasets.
@@
Additionally, we analyze the global token embedding distribution in LLMs and pre-trained vision models using datasets included in the models’ training corpus (Touvron et al., 2023; Grattafiori et al., 2024; Sun et al., 2024b). Figure 2 plots token input frequency distributions and their occurrences in the dataset on a log scale, revealing a scale-free structure among the token embeddings that follow power-law decay. The figure also shows token norm distributions for learned embeddings, where the count for high-norm embeddings decreases exponentially at the right tail, reinforcing the scale-free property. This scale-free organization suggests an underlying hierarchical structure (Barabási & Albert, 1999), where a small number of high-frequency tokens act as hubs within the semantic network. Such hierarchies are naturally represented in hyperbolic embedding spaces (Krioukov et al., 2010; Nickel & Kiela, 2017), which efficiently encode tree-like structures due to their exponential volume growth. The non-Euclidean structures in token distribution are exhibited to some extent even in Euclidean models and are most likely attributed to the models being optimized during training to maximize representational quality. However, the scale-free properties are still not yet fully captured by the Euclidean foundational model, where the count of embeddings with small norms still increases. See Appendix B for statistics of more models and the details of the datasets. We do observe that the non-Euclidean structure in the final embeddings is significantly reduced compared to earlier layers, even though some non-Euclidean structure persists in the final embeddings. Nevertheless, this does not diminish the presence of non-Euclidean structures in the model’s handling of inputs through token embeddings. After all, final embeddings focus on supporting prediction tasks after intermediate Euclidean transformations such as attention layers and layer normalization.
@@
Additional Structures. In addition to hierarchical structure, data may exhibit other structural characteristics, such as cycles and loops. Foundation models built on Euclidean geometry struggle to represent such structures without significant distortion. Many real-world tasks, such as 3D shape analysis (Esteves et al., 2017, 2019), medical imaging (Bekkers et al., 2018; Winkels & Cohen, 2018), and physics-informed machine learning (Liao & Smidt, 2023; Liao et al., 2024; Anderson et al., 2019; Cohen et al., 2018), can benefit from encoding data geometry as inductive bias. Euclidean operations, such as the convolutional layers, encode only translation invariance (Esteves et al., 2020), resulting in performance limitations for these tasks.
@@ 3.3 The Necessity of Non-Euclidean Geometry for Foundation Models
Building on the previous findings, we further explore how encoding suitable geometric structure can improve foundation model performance.
@@
(1) Addressing the limitations of Euclidean space in capturing intrinsic token structures. Recent research shows that the attention mechanism plays a pivotal role in the expressive capacity of LLMs (Aghajanyan et al., 2021; Song & Zhong, 2023; Vaswani et al., 2017; Balestriero et al., ).
Lemma 3.4 (Causal Single-Head Convex Hull (Balestriero et al., )).
@@
Let $X\in\mathbb{R}^{T\times D^{(\ell)}}$ be the input to the $\ell$-th layer of an LLM, where $T$ is the sequence length and $D^{(\ell)}$ is the feature dimension. The output of attention head $h$ at position $i$ lies within the convex hull of the first $i$ rows of $XV_{h}^{(\ell)}$: $\text{Head}_{h}^{(\ell)}(X)_{i}\in\text{Hull}\left\{(V_{h}^{(\ell)})^{\top}x_{j}\mid j=1,\dots,i\right\}.$ Its effective dimension is bounded by the number of nonzero attention weights: $\text{dim}_{\text{eff}}\leq\#\left\{\text{Attn}_{h}^{(\ell)}(X)_{i,j}>0\mid j\in\{1,\dots,i\}\right\}.$ Here, $\text{Attn}_{h}^{(\ell)}(X)$ is the attention matrix for head $h$ at layer $\ell$, computed as: $\text{Attn}_{h}^{(\ell)}(X)=\text{softmax}_{\text{causal}}(XQ_{h}^{(\ell)}K_{h}^{(\ell)\top}X^{\top}).$
@@
This lemma highlights that next-token prediction in LLMs is strongly influenced by relationships encoded in previous tokens. Standard attention mechanisms assume an Euclidean representation space, where token embeddings and their pairwise relationships fit this geometry. However, as investigated in Table 2, each prompt exhibits non-Euclidean characteristics. Consequently, the Euclidean attention mechanism fails to model hierarchical syntax, circular or periodic dependencies, and other complex relational structures without distortion, as demonstrated in Section 3.1. Replacing the Euclidean attention matrix with a non-Euclidean one (e.g., hyperbolic or spherical) could better capture previous token relationships, aligning with the intrinsic data structure and enhancing next-token prediction. For example, hyperbolic geometry compresses distances exponentially, ensuring that distant but structurally related tokens (e.g., a root concept and its distant occurrences in a prompt / document) remain meaningfully close. This enables attention mechanisms to efficiently preserve hierarchical relationships and capture long-range dependencies.
@@
(2) Leveraging distortion-dimension trade-offs in hierarchical concept modeling. Recent studies extend beyond representation space analysis to unify it with the unembedding space, examining how Euclidean-based LLMs encode hierarchies geometrically (Park et al., 2024a, b). An LLM consists of a mapping function $\lambda$ and an unembedding layer. The function $\lambda$ maps input text $x$ to a vector $\lambda(x)\in\mathbb{R}^{d}$, while the unembedding layer assigns $\gamma(y)\in\mathbb{R}^{d}$ to each token $y$. The token probability distribution follows: $P(y\mid x)=\frac{\exp(\lambda(x)^{\top}\gamma(y))}{\sum_{y^{\prime}\in\text{Vocab}}\exp(\lambda(x)^{\top}\gamma(y^{\prime}))}.$ To unify the different spaces, the embedding and unembedding spaces can be reformulated using transformations $g(y)=A(\gamma(y)-\bar{\gamma}_{0}),\quad\ell(x)=A^{-\top}\lambda(x),$ where the Euclidean inner product serves as the causal inner product.
This framework, validated through experiments, reveals that Euclidean-based LLMs encode hierarchical concepts orthogonally, modeling a tree structure where parent (e.g., animal) and child (e.g., bird, mammal) vectors are perpendicular. However, this orthogonal representation demands high dimensionality, which scales significantly as hierarchical relationships grow more complex (Gupta, 1999). Non-Euclidean spaces offer a more efficient alternative, preserving hierarchical relationships while significantly reducing dimensionality (Nickel & Kiela, 2017, 2018).
@@
(3) Surpassing Euclidean constraints in modeling multi-modal heterogeneity. Data from different modalities exhibit significant distributional variations, influenced by contextual factors, use cases, cultural differences, and varying interpretations of the same information. This complexity intensifies in multi-modal data, where each modality (e.g., images, text, audio) has distinct hierarchical structures (Li et al., 2022; Wei et al., 2023; Liang et al., 2022; Goel et al., 2022; Jiang et al., 2023). Many studies highlight this issue. Liang et al. (2022) showed that the latent modality gap and distinct modality structures emerge in the latent space due to initialization and the contrastive learning process, impacting downstream tasks. Wang et al. found that different modalities exist on separate manifolds, making a unified Euclidean foundation model highly redundant in parameters and requiring varying degrees of pruning for different modalities. This suggests that a unified Euclidean geometry struggles to capture multi-modal, cross-domain relationships, as its flat structure lacks the expressiveness needed for complex, multi-faceted interactions.
@@
Hyperbolic geometry excels in vision-language foundation models by effectively capturing hierarchical relationships, enhancing representation interpretability (Desai et al., 2023; Pal et al., 2024; Ramasinghe et al., 2024). Its advantages extend to multimodal tasks, including image-video-skeleton (Li et al., 2024) and video-audio applications (Hong et al., 2023). For example, in hyperbolic embeddings, higher-level hierarchical concepts lie closer to the origin with more specific concepts residing in more peripheral regions, enabling geodesic reasoning when navigating through concept hierarchies. Hyperbolic embeddings minimize distortion while preserving real-world data structures, reinforcing the necessity of non-Euclidean geometry in foundation models. More broadly, generalized non-Euclidean geometries provide diverse curvatures, from positive to negative, enabling multiple manifolds to encode distinct data distributions (Wilson et al., 2014; Gao et al., 2022, 2021), surpassing the constraints of flat Euclidean space.
In summary, we argue that moving beyond Euclidean geometry is essential for next-generation foundation models. Euclidean models struggle with capturing token structures, face dimensionality limits, and fail to model multimodal heterogeneity, requiring high dimensions but lacking geometric adaptability. Non-Euclidean geometry, with its flexible curvature and manifold structures, overcomes these limitations in distortion, scalability, and adaptability, providing a more scalable and expressive framework for foundation models.
@@
Figure 3: The roadmap for integrating non-Euclidean geometries into foundation models, includes (a) fine-tuning existing Euclidean foundation models, (b) pretraining from scratch, and (c) hybrid architectures. Four strategies are shown in (a), labeled with circled numbers 1-4, respectively: geometric prompt tuning, geometric low-rank adaptation, geometric knowledge distillation, and geometric transfer learning. All learnable components are highlighted in red in (a) and (c). We only show one decoder for each strategy for simplicity.

### v2

@@ 3.1 Limitations of Euclidean Geometry for Foundation Models
[robinson2006sphere] Let $(X,d_{X})$ be a metric space with $X=\{a,b,c,d\}$ and $d_{X}(a,b)=d_{X}(a,c)=d_{X}(a,d)=2L$ and $d_{X}(b,d)=d_{X}(c,d)=L$ for $L\in\mathbb{R}^{+}$. Then $X$ admits no isometric embedding into $\mathbb{R}^{n}$ for any $n$.
As these points can be isometrically embedded into $S^{k}$, Theorem˜3.3 shows that $S^{k}$ cannot be isometrically embedded into $\mathbb{R}^{n}$ for any $n\in\mathbb{N}$, resulting in distortion when encoding rotational equivariance. In contrast, non-Euclidean geometry can provide a more natural representation of complex topological structures, reducing distortion and dimensionality of the embedding space. For instance, [sarkar2011low] showed that every finite tree admits an embedding into the hyperbolic plane $\mathbb{H}^{2}$ with $1+\epsilon$ multiplicative distortion for any $\epsilon>0$, leading to $O(1)$ distortion with low dimensionality.
Take-away. The implications of the previous theoretical discussion are numerous: (1) Limited scalability. Theorem˜3.1 highlights the distortion-dimension trade-off for Euclidean foundation models when embedding complex structures, which is reflected in the computational resources required in these models. Non-Euclidean geometry produces higher quality embeddings in significantly lower dimensions, offering enhanced model scalability; (2) Performance bottleneck. Theorem˜3.2 and 3.3 demonstrate that even in the case of an abundance of compute resources, the linear assumption in Euclidean foundation models could still incur significant distortion regardless of the embedding dimension for a wide range of topological structures, resulting in a performance upper bound. These theoretical results are validated in Section˜3.2.
Model arXiv C4 Common Crawl GitHub StackExchange Wikipedia
RoBERTa-Base [liu2019roberta] $0.15\pm 0.06$ $0.18\pm 0.04$ $0.17\pm 0.04$ $0.12\pm 0.04$ $17\pm 0.07 0.0$$07\pm 0.05$
LLaMA3.1-8B [dubey2024llama3] $0.15\pm 0.05$ $0.16\pm 0.07$ $0.15\pm 0.06$ $0.12\pm 0.05$ $0.18\pm 0.06$ $0.10\pm 0.04$
GPT-NeoX-20B [gpt-neox-20b] $0.14\pm 0.03$ $0.17\pm 0.06$ $5\pm 0.05 0.11±0$$1\pm 0.04$ $14\pm 0.04 0.0$$09\pm 0.03$
Gemma2-9B [gemma2] $0.17\pm 0.06$ $0.19\pm 0.04$ $0.20\pm 0.05$ $0.15\pm 0.05$ $0.18\pm 0.04$ $0.15\pm 0.03$
Reference $\delta$ values $0.99\pm 0.01$ $0.63\pm 0.01$ $0.40\pm 0.04$ $14\pm 0.01 0.0$0.0 -
Figure 2: Token frequency v.s. token count (left 2) and token norm vs token count (right 2) for LLaMa3.1-8B and LLaMaGen. The datasets are chosen to be within the training corpus. The token-frequency figures show the scale-free properties of the token inputs. The token norm figures reflect this property for learned token embeddings to some extent, with token count increasing exponentially for high-normed tokens at the left tail. However, the Euclidean embeddings still do not fully capture this property and deviate from it at the right tail. More statistics are shown in Appendix˜B.
3.2 Non-Euclidean Geometry in Foundation Models
In this section, we empirically assess embedding distortions for different geometries to validate our claims in Section˜3.1 and demonstrate that non-Euclidean geometry is more suitable. We then analyze token embeddings in foundation models, showing that structures that align with non-Euclidean geometry are prevalent, highlighting the need for alternative geometric frameworks.
Table 2: Average (point-wise) distortion on canonical graphs with $96$ nodes, comparing four spaces with total dimension 66. The least distortion is achieved by the space with the most suitable geometry.
@@ 3.2 Non-Euclidean Geometry in Foundation Models
Empirical Validation. We empirically validate our claim that Euclidean space fails to capture complex structures faithfully and that non-Euclidean spaces are better suited for producing high-quality embeddings. Table 2 compares the average (point-wise) distortion of four geometric spaces ($\mathbb{R}^{6}$, $\mathbb{H}^{-1,6}$, $\mathbb{S}^{1,6}$, and $\mathbb{H}^{-1,3}\times\mathbb{S}^{1,3}$) in representing three canonical graphs (Tree, Cycle, and Ring of Trees) with 96 nodes, each corresponding to a different type of intrinsic token relationships (hierarchical, cyclical, and both). We intentionally dismissed invariance from kernel designs—both Euclidean and non-Euclidean, e.g. orthogonal invariance (see Section˜3.3)—to isolate the representational capacity of the spaces. The most suitable geometry varies by graph type—Lorentzian space ($\mathbb{H}^{-1,6}$) for trees, spherical space ($\mathbb{S}^{1,6}$) for cycles, and mixed geometry ($\mathbb{H}^{-1,3}\times\mathbb{S}^{1,3}$) for rings of trees—emphasizing the importance of selecting an appropriate geometry to minimize distortion. As a fundamental desired property of foundation models is for them to generalize to a wide variety of data, non-Euclidean geometry enables significantly increased flexibility in the choice of geometry for embedding spaces, such as the use of a product of manifolds of varying curvatures (positive, zero, and negative) that encompasses the Euclidean space.
Figure 3: Distortion for embedding a Tree with 96 nodes for varying dimensionality (log scale). Non-Euclidean geometry achieves smaller distortion with significantly fewer dimensions and has better scaling.
We also compute the distortion value against varying dimensionality. An example is shown in Figure˜3 for the case of a tree with 96 nodes, plotted on log-scale for visibility. The hybrid manifold is a product of hyperbolic and spherical spaces, each with half the dimension. The 4-dimensional hyperbolic space achieves a significantly smaller distortion than Euclidean embeddings with 50 dimensions. This reflects Takeaway 1 in Section˜3.1: non-Euclidean geometry achieves superior performance with significantly fewer dimensions. Distortion also continues to decrease for hyperbolic and hybrid spaces but plateaus for Euclidean space, reflecting Takeaway 2, where Euclidean space has theoretical upper bounds for embedding trees but non-Euclidean geometry has the potential to continue the performance scaling law at high dimensionality. See Appendix˜B for additional plots of other graph types.
@@
Hierarchies in Token Embeddings. Based on the above results validating Euclidean embeddings’ limitations, we further show that these structures which Euclidean spaces struggles to embed is prevalent in foundation models. To explore the intrinsic structure within the representations of foundation models, we utilize $\delta$-hyperbolicity [gromov1987hyperbolic], which quantifies the extent to which a metric space deviates globally from a tree metric (see Appendix˜C). Each token is treated as a point in a discrete metric space $X$, and a graph is constructed based on similarity scores between each pair. We evaluate the hyperbolicity of token embeddings in LLMs, where lower values suggest a tree-like structure. As shown in Table˜1, the consistently low $\delta$-hyperbolicity values suggest hierarchical structures within each prompt across diverse datasets.
@@
We also analyze the global token embedding distribution in LLMs and pre-trained vision models using datasets included in the models’ training corpus [touvron2023llama, dubey2024llama3, sun2024autoregressive]. Figure˜2 plots token input frequency distributions and their occurrences in the dataset on a log scale, revealing a scale-free structure among the token embeddings. This scale-free organization suggests an underlying hierarchical structure [barabasi1999emergence], where a small number of high-frequency tokens act as hubs within the semantic network. The figure also shows token norm distributions for learned embeddings, where the count for high-norm embeddings increases exponentially at the left tail, reinforcing the scale-free property. The non-Euclidean structures in token distribution are exhibited to some extent even in Euclidean models are most likely attributed to the models being optimized during training to maximize representational quality. However, the scale-free properties are still not yet fully captured by the Euclidean foundational model, where the count of embeddings with large norms still decreases at the right tail. See Appendix˜B for more statistics.
Token Hierarchies to Model Performance. While limited works have explored how embedding distortions connect to model performance and concrete connection has not been established due to a variety of challenges (see Section˜5), recent works have made preliminary efforts toward this goal. Recent efforts have shown that hyperbolic LLMs outperform Euclidean counterparts at multi-choice question answering tasks while exhibiting better semantic hierarchy modeling and better separation of words based on specificity, where more specific words are embedded further away from the origin [he2025helm]. These results suggest that the failure of Euclidean geometry to model complex structure is a potential cause of degraded downstream performance.
@@
Additional Structures. In addition to hierarchical structure, data may exhibit other structural characteristics, such as cycles and loops. Many real-world tasks, such as 3D shape analysis [esteves2017learning, esteves2019equaivariant], medical imaging [bekkers2018roto, winkels20183d], and physics-informed machine learning [liao2023equiformer, equiformer_v2, anderson2019cormorant, cohen2018spherical], can benefit from encoding data geometry as inductive bias. Euclidean operations, such as convolutional layers, encode only translation invariance [esteves2020swscnn], resulting in performance limitations for these tasks.
@@ 3.3 The Necessity of Non-Euclidean Geometry for Foundation Models
Here we further explore how non-Euclidean geometry could improve foundation model performance.
@@
(1) Addressing the limitations in capturing intrinsic token structures. Recent research shows that the attention mechanism plays a pivotal role in the expressive capacity of LLMs [aghajanyan2021intrinsic, song2023uncovering, vaswani2017attention, balestrierocharacterizing].
Lemma 3.4 (balestrierocharacterizing).
@@
Let $X\in\mathbb{R}^{T\times D_{(\ell)}}$ be the input to the $\ell$-th layer of an LLM, where $T$ is sequence length and $D_{(\ell)}$ is feature dimension. Attention head $h$’s output at position $i$ is in the convex hull of the first $i$ rows of $XV_{h,{(\ell)}}$: $\text{Head}_{h,{(\ell)}}(X)_{i}\in\text{Hull}\left\{(V_{h,{(\ell)}})^{\top}x_{j}\mid j=1,\dots,i\right\}.$ with bounded effective dimension: $\text{dim}_{\text{eff}}\leq\#\left\{\text{Attn}_{h,{(\ell)}}(X)_{i,j}>0\mid j\in\{1,\dots,i\}\right\}.$ Here, $\text{Attn}_{h,{(\ell)}}(X)$ is the attention matrix for head $h$ at layer $\ell$: $\text{Attn}_{h,{(\ell)}}(X)=\text{softmax}_{\text{causal}}(XQ_{h,{(\ell)}}K_{h,{(\ell)}}^{\top}X^{\top}).$
@@
This lemma highlights that next-token prediction in LLMs is strongly influenced by relationships encoded in previous tokens. As shown in Table 1, tokens exhibits non-Euclidean characteristics. Consequently, the standard Euclidean attention mechanism does not faithfully capture hierarchical syntax, periodic dependencies, and other complex token relationships, as demonstrated in Section˜3.1. Utilizing non-Euclidean attention mechanisms instead could better capture previous token relationships by aligning with the intrinsic data structure, thus enhancing next-token prediction. For example, hyperbolic geometry compresses distances exponentially, ensuring that distant but structurally related tokens (e.g., a root concept and its distant co-occurrences in a prompt) remain meaningfully close, enabling attention mechanisms to efficiently capture long-range dependencies and hierarchies.
@@
(2) Alleviating distortion-dimension trade-offs. Recent studies examined how Euclidean-based LLMs encode hierarchies geometrically [park2024geometry, parklinear], where a mapping function $\lambda$ maps input text $x$ to a vector $\lambda(x)\in\mathbb{R}^{d}$, and an un-embedding layer assigns $\gamma(y)\in\mathbb{R}^{d}$ to each token $y$. The token probability distribution is given by $P(y\mid x)=\frac{\exp(\lambda(x)^{\top}\gamma(y))}{\sum_{y^{\prime}\in\text{Vocab}}\exp(\lambda(x)^{\top}\gamma(y^{\prime}))}.$ To unify the different spaces, the embedding and unembedding spaces can be reformulated using transformations $g(y)=A(\gamma(y)-\bar{\gamma}_{0}),\quad\ell(x)=A^{-\top}\lambda(x),$ where the Euclidean inner product serves as the causal inner product. This framework shows that Euclidean LLMs encode hierarchical concepts orthogonally, where parent (e.g., animal) and child (e.g., bird, mammal) vectors are perpendicular. Yet, since provides only d orthogonal dimensions, Euclidean spaces must rely on high dimensionality to capture the expansive semantic hierarchies of language [gupta1999embedding]. Non-Euclidean spaces offer a more efficient alternative, preserving hierarchical relationships while significantly reducing dimensionality [nickel2017poincare, nickel2018learning].
@@
(3) Improved multi-modal heterogeneity modeling. Data from different modalities vary significantly due to contextual factors, use cases, cultural differences, and different interpretations of the same information. This complexity intensifies in multi-modal data, where each modality has distinct complex structures [li2022clmlf, wei2023tackling, liang2022mind, goel2022cyclip, jiang2023understanding]. For instance, latent modality gap and distinct modality structures exist in the latent space due to initialization and the contrastive learning process, impacting downstream tasks [liang2022mind]. Different modalities also exist on separate manifolds [wangexploring], making a unified Euclidean foundation model highly redundant in parameters and requiring varying degrees of pruning for different modalities. Thus, Euclidean space struggles to capture multi-modal cross-domain relationships, as its flat structure lacks the flexibility needed for multi-faceted interactions in the data.
@@
Non-Euclidean spaces exhibit much more geometric flexibility to enable multiple manifolds that encode different data distributions [wilson2014spherical, gao2022curvature, gao2021curvature]. For instance, hyperbolic geometry excels in vision-language foundation models by effectively capturing hierarchical relationships [desai2023hyperbolic, pal2024compositional, ramasinghe2024accept], improving performance in tasks such as image-video-skeleton [li2024isolated] and video-audio applications [hong2023hyperbolic] while enhancing representation interpretability—higher-level hierarchical concepts lie closer to the origin with more specific concepts residing in more peripheral regions, enabling geodesic reasoning when navigating through concept hierarchies.
@@
Figure 4: Roadmap for integrating non-Euclidean geometries into foundation models, includes (a) fine-tuning existing Euclidean foundation models, (b) pretraining from scratch, and (c) hybrid architectures. Four strategies are shown in (a), labeled with circled numbers 1-4, respectively: geometric prompt tuning, geometric low-rank adaptation, geometric knowledge distillation, and geometric transfer learning. All learnable components are highlighted in red in (a) and (c).